OpenStudy (anonymous):

When doing homework, a student uses the multiplication property of equality to solve for c below. Which statement BEST applies to the sample? 5c = 10 5c . 1/5 = 10 . 1/5 c=2 A. The student failed to keep both sides of the equation balanced. B. The student’s final answer does not have the variable c isolated. C. The use of the multiplication property of equality is sound, and the answer is correct. D. The student should have used the division property of equality.

OpenStudy (yellowlegoguy99):

5c = 10 5c * 1/5 = 10 * 1/5 c = 2

OpenStudy (yellowlegoguy99):

The answer is C. Medal?

OpenStudy (jojo4eva):

@yellowlegoguy99 if he didnt say he was going to give you one if you answered the question right then dont expect one. You shouldnt juss be answering questions just to get a medal you should be helping so that people will be assisted and will have the best understanding as possible of what they are asking so stop demanding people to give you a medal